# Do Construction Distributions Shape Formal Language Learning In German BabyLMs?

### v1

@@ Abstract
We analyze the influence of utterance-level construction distributions in German child-directed speech on the resulting formal linguistic competence and the underlying learning trajectories for small language models trained on a novel collection of developmentally plausible language data for German. We find that trajectories are surprisingly robust for markedly different distributions of constructions in the training data, which have little effect on final accuracies and almost no effect on global learning trajectories. While syntax learning benefits from more complex utterances, lexical learning culminates in better scores with more fragmentary data. We argue that LMs trained on developmentally plausible data can contribute to debates on how rich or impoverished linguistic stimuli actually are.
@@ 1 Introduction
One of the most contentious issues in language acquisition is the relationship between the input that the learner receives and the resulting linguistic system. Child-directed speech (or CDS) is structurally simple and, especially in the first three years of life, abounds with questions, imperatives, and fragmentary utterances, but features very few complex sentences Cameron-Faulkner et al. (2003); Bunzeck and Diessel (2024). This distribution of utterance-level constructions is conducive to the functional side of language acquisition: caregivers talk in this way to elicit responses, steer behavior or establish joined attention. But how do children acquire full-fledged, formal grammatical knowledge from such supposedly skewed input? Generativist approaches see this kind of stimulus as too impoverished to kickstart formal linguistic development and assume an underlying, innate grammatical endowment that enables it (cf. Chomsky, 1965; Crain and Pietroski, 2001; Thomas, 2002; Berwick et al., 2011), whereas usage-based scholars argue that domain-general learning mechanisms are sufficient to develop a full-fledged mental linguistic system from this input, which is rich enough (cf. MacWhinney, 2004; Tomasello, 2005).
The connectionist “renaissance”, fueled by the increased adoption of deep learning and transformer language models, has opened up new avenues of addressing such questions Warstadt and Bowman (2022). Developmental AI Dupoux (2018) and the BabyLM community Warstadt et al. (2023); Hu et al. (2024); Charpentier et al. (2025) have demonstrated that supposedly tabula rasa learners can indeed acquire syntactic structures, the formal side of language, successfully from very little data Huebner et al. (2021), but without tapping into the functional side of language (Mahowald et al., 2024). This makes them ideal testbeds for the aforementioned issue: is the construction distribution found in child-directed speech really too impoverished to learn formal linguistic capabilities, and are other, more complex registers, e.g. book text Cameron-Faulkner and Noble (2013); Noble et al. (2018), a richer and more beneficial form of input?
@@
To investigate this, we compile a novel German BabyLM training set, for which we conduct the first utterance-level construction analysis for German. We find that distributions align with findings for English and other langauges, and that child-directed media (e.g., children’s encyclopedias) provide a construction-level middle ground between CDS and more complex book text. We then create three 5M-token subsets with distinct constructional profiles, e.g., varying the proportion of fragmentary and complex utterances, and train small, character- and subword-level Llama models on them. Our evaluation with lexical, syntactic, and semantic minimal pairs Bunzeck et al. (2025); Mueller et al. (2020); He et al. (2025) reveals that differences between grammatically complex training data and a developmentally plausible constructional distribution are fairly small. While certain syntactic phenomena are learned somewhat better from more complex sentences, lexical learning improves with more fragments and questions in the input. Most interestingly, input complexity only modulates the steepness of the resulting learning trajectories, but has no principal effect on the amount of input needed to kickstart learning.
@@ 2 Constructions in children’s input
Child-directed speech can be seen as a separate linguistic register and is the primary input that children encounter in their first years. It has mostly been scrutinized from phonetic and lexical viewpoints, where well-established findings are that it features slower speech and exaggerated intonation patterns which infants prefer listening to Zangl and Mills (2007), while its vocabulary is rather restricted to everyday topics and children’s immediate surroundings Snow and Ferguson (1977). Structurally, child-directed utterances are usually shorter and simpler than adult-directed ones Genovese et al. (2020) and feature high amounts of structural and lexical repetition Tal et al. (2024). Statistical properties of the input directly influence the children’s order of acquisition for syntactic patterns Huttenlocher et al. (2002); Ambridge et al. (2015), e.g. for relative clauses Diessel and Tomasello (2000); Brandt et al. (2008); Chen and Shirai (2015).
@@
Early studies were mostly concerned with mapping out how much CDS is ungrammatical or otherwise “wrong” (in the sense of hesitations, false starts, etc., cf. Pine, 1994), but the quantitative turn in linguistics Janda (2013) has enabled more holistic analyses. In a seminal study, Cameron-Faulkner et al. (2003) analyze utterance-level constructions in child-directed English via a large corpus of toyplay sessions featuring children and their caregivers. They show that CDS features only few “canonical” SV(X)-utterances but abounds with questions, lexical fragments, or copula constructions. The reported construction distributions also hold for typologically different languages, e.g., Irish Cameron-Faulkner and Hickey (2011). These constructions and their real-world functions help children to quickly understand the functional side of language. On the other hand, the most common and repetitive utterances that English-speaking children hear from their mothers represent a rather skewed sample of the presumed, underlying formal language system. Generativist approaches would argue that certain formal structures, e.g. question formation from relative clauses, are not attainable from this kind of language, as the input never contains specific examples Chomsky (1980) (although Pullum and Scholz (2002) argue that the input frequently contains exactly such specific examples). Conversely, constructivist approaches, which view language learning as the re-construction of the target language Behrens (2021), argue that this kind of input is actually conducive to formal aspects of acquisition, by providing anchor points for first words and their semantic links to real-world reference, which then serve as building blocks for a gradual development into larger schemas (like questions with relative clauses).
@@
Although CDS features such a skewed construction distribution, written language aimed at children, i.e., in the form of children’s books, is characterized by a much higher rate of canonical SV(X)-constructions than found in CDS Cameron-Faulkner and Noble (2013). In contrast, questions rarely occur in books. CDS produced in shared book reading presents a middle-ground — it contains more complex and SV(X)-constructions than regular CDS, but less than book text alone Noble et al. (2018). They argue that shared reading therefore plays an important role in moving children from early, isolated traces of linguistic knowledge to a rich mental language system. This also aligns with the findings by Bunzeck and Diessel (2024), who show that the distribution of constructions in CDS varies with situation type (toyplay features most questions, meal sessions beget more imperatives, shared book reading features more complex constructions) and child age (questions and imperatives become less frequent with age). They suggest that CDS is therefore adapted to support children’s cognitive and linguistic development. Yet, as corpus studies are necessarily descriptive and cannot establish causal/mechanistic connections on their own (e.g. what would happen if a child never hears CDS), it remains questionable if this is actually true.
@@ 3 Input in developmentally plausible LMs
Early approaches to modeling language acquisition with neural networks used hand-picked, manually ordered data points Rumelhart and McClelland (1986) or synthetic data generated with hand-crafted grammars Elman (1993); Christiansen and Chater (1999); Chang et al. (2006). Both lack developmental plausibility. Since then, data availability has improved with the establishment of developmental corpora. Frequently, CDS from the CHILDES database MacWhinney (2000) is used to train developmentally plausible LMs (cf. Pannitto and Herbelot, 2020; Huebner et al., 2021). While CHILDES-based models have the advantage of learning from authentic data only, they have the disadvantage of not accessing the full breadth of the linguistic input children receive. Children are exposed to many more different registers of language throughout their linguistic development, like the aforementioned shared (or solitary) book reading, or television shows Montag (2019); Gowenlock et al. (2024). As a response to this, the BabyLM corpora propose a data mix of varied spoken and written sources, from CDS over adult-adult conversations to OpenSubtitles Lison and Tiedemann (2016), but also children’s Hill et al. (2015) and adults’ books Gerlach and Font-Clos (2020). All data included in them could be plausibly encountered by children, which provides opportunities to ablate the influence of architecture/training on the learned linguistic knowledge.
For languages other than English, data availability is the greatest problem for the composition of developmentally plausible datasets. Salhan et al. (2024) use only data available from CHILDES for French, German, Japanese, and Chinese models, whereas Prévot et al. (2024) compare models trained on spoken data (child-directed + adult-adult conversations) with models trained on the French Wikipedia. As such, these first forays into more polyglot BabyLMs are still constrained to the child-directed input found in CHILDES and do not extend to the variety of inputs that children are exposed to Soderstrom (2007); Gowenlock et al. (2024). Notably, Suozzi et al. (2025) introduce an Italian BabyLM but do not elaborate on their data sources beyond CHILDES.
Linguistic properties
The actual linguistic make-up of pre-training corpora and its influence on linguistic performance have only recently begun to receive increased scrutiny. Focusing on the lexical level, Yam and Paek (2024) measure sentence-level textual complexity with readability metrics based on text-wide word/syllable–sentence ratios for different corpora (CHILDES only, BabyLM corpus, synthetic data, Project Gutenberg). They find that models trained on more complex text perform better at syntactic benchmarks, but simpler data (CHILDES) is learned better in terms of perplexity and loss convergence. Also on the lexical level, Muckatira et al. (2024) filter regular, non-BabyLM pre-training corpora for text spans that only contain vocabulary also found in English CHILDES data and find that simplified models generate more coherent text than models trained on more complex data and also succeed in syntactic tests if the test data is filtered accordingly. In contrast, Edman et al. (2024) change the semantic content of the pre-training data and use datasets that approximate linguistic input that second-language learners get, e.g., dictionary entries, grammar books, and paraphrased sentences. While grammar books moderately improve syntactic evaluation, there is no positive effect observable for the addition of the other text types.
@@
While actual research on the structural/syntactic properties of the input is still rather sparse, training on filtered corpora has been used in pilot studies. Patil et al. (2024) and Misra and Mahowald (2024) filter out specific grammatical constructions from the standard BabyLM corpora and then probe the resulting models for knowledge of these grammatical constructions (which might also be analogically learned from related constructions or constructed from their parts). Patil et al. (2024) show that their models succeed on the BLiMP benchmark Warstadt et al. (2020), even if sentences containing structures targeted in BLiMP’s minimal pair sets are removed. Similarly, Misra and Mahowald (2024) show that acceptability scores for the English AANN construction can be reliably estimated from models that have never seen this specific construction. In sum, then, models appear to be able to generalize from indirect evidence and learn language in a somewhat constructivist, bottom-up fashion.
So far, previous work in NLP has failed to adequately investigate the structural composition of child-directed data. Most studies focus on lexical or semantic properties, emphasizing content over structure, and child-directed data is usually equated with a somewhat fitting vocabulary or with just being authentic data. However, findings from usage-based linguistics suggest that structural properties, like utterance-level construction distributions, play a crucial role in language acquisition. Patil et al. (2024) and Misra and Mahowald (2024) remove specific constructions from their data, whereas we aim to explore whether different global distributions of constructions influence the resulting linguistic knowledge and learning trajectories.
@@ 4 A German BabyLM dataset
To construct a German dataset, we use a variety of developmentally plausible sources, similar to the English BabyLM data Warstadt et al. (2023); Choshen et al. (2024). We use i) all data from German CHILDES corpora MacWhinney (2000), including frog stories from TalkBank Berman and Slobin (1994) and math lessons from ClassBank Stigler et al. (2000), ii) subtitles from OpenSubtitles Lison and Tiedemann (2016), iii) adult conversations from the CallHome corpus Karins et al. (1997), and iv) written data from Project Gutenberg, from which we downloaded a manually curated sample of children’s books, young adult literature and literature commonly read in German schools. We supply this data with two corpora, the DreamBank database of self-reported dreams Domhoff and Schneider (2008) and short news texts from the Leipzig corpus Goldhahn et al. (2012), although they are not child-directed per se, these sources are child-available in everyday language.
To approximate child-available input even better, we tap into freely available child/learner-directed sources and compile four additional subcorpora for our dataset. The Wikibooks Wikijunior shelve features educational resources aimed at children, focusing on a diverse array of topics such as technology or nature. The Klexikon is a children’s wiki in German, featuring more than 3,000 articles aimed at children between 5–15. A simplified version of it is the MiniKlexikon, which features over 1,500 articles aimed at beginning readers. Finally, we also scrape the complete archives of Fluter, a magazine aimed at young adults published by the Federal Agency for Civic Education, which contains a large body of non-fiction. All resources are CC-licensed. Table 1 shows the raw token numbers for all corpora. In sum, we get 16.5M lexical tokens. Before further analysis, we extensively clean and normalize our data (details in Appendix C). We share our dataset on huggingface.

### v2

@@ Abstract
We analyze the influence of utterance-level construction distributions in German child-directed/child-available speech on the resulting word-level, syntactic and semantic competence (and their underlying learning trajectories) in small LMs, which we train on a novel collection of developmentally plausible language data for German. We find that trajectories are surprisingly robust for markedly different distributions of constructions in the training data, which have little effect on final accuracies and almost no effect on global learning trajectories. While syntax learning benefits from more complex utterances, word-level learning culminates in better scores with more fragmentary utterances. We argue that LMs trained on developmentally plausible data can contribute to debates on how conducive different kinds of linguistic stimuli are to language learning.
@@ 1 Introduction
One of the most contentious issues in language acquisition is the relationship between the input that learners receive and the resulting linguistic system Pullum and Scholz (2002); Clark and Lappin (2011). Child-directed speech (or CDS) is structurally simple: Especially in the first three years of life, it abounds with questions, imperatives, and fragmentary utterances, but features fewer SV(X) and very few complex sentences, which instantiate “canonical” word order Cameron-Faulkner et al. (2003). This distribution of utterance-level constructions is conducive to the functional side of language acquisition: caregivers talk in this way to elicit responses, steer behavior, or establish joint attention. But how do children acquire full-fledged, formal grammatical knowledge from such supposedly skewed input? While its advantages for aspects like speech segmentation or word learning are somewhat accepted Yurovsky et al. (2012); Cristia et al. (2019), its influence on syntax remains debated: whereas some generativist approaches see any kind of input as too impoverished to learn a full-fledged syntactic system (cf. Chomsky, 1965; Crain and Pietroski, 2001; Guasti, 2002; Thomas, 2002; Berwick et al., 2011), constructionist and usage-based scholars argue that this supposedly skewed input actually aids syntax learning MacWhinney (2004); Tomasello (2005); Bunzeck and Diessel (2024).
Project Gutenberg: Complex sentences Aber sie war in Angst, dass wir die Larven beschädigen würden, die zu Arbeiterinnen heranwachsen sollten. (But she was afraid that we would damage the larvae which were supposed to grow into workers.)
MiniKlexikon: Transitive SP sentences Der Grafiker entwirft das Bild vorne auf dem Buch. (The graphic designer designs the picture on the cover of the book.) Der Friseur schneidet die Haare. (The hairdresser cuts the hair.)
Child-directed speech: Fragmentary utterances noch mehr! (even more!) ja. (yes.) mit dem Flugzeug. (with the airplane.)
Figure 1: Examples for most frequent construction types from different portions of our German BabyLM corpus
The connectionist “renaissance”, fueled by deep learning and Transformer language models, has opened up new avenues of investigating the relationship between an artificial learner’s acquired linguistic system and the nature of its training data, more recently also from a constructionist/non-generativist viewpoint Weissweiler et al. (2023); Piantadosi (2024). LLMs, pretrained on raw language data only, and instruction-finetuned chatbots based on them, generate text without grammatical errors, and perform well in controlled syntactic test suites. Unfortunately, though, their massive parameter size does not preclude the possibility that their linguistic capabilities result from memorization rather than generalization Millière (2024). Furthermore, the sheer amount of their pretraining data exceeds human learner’s input by many orders of magnitude, putting their relevance for linguistic modeling into question. Work within the BabyLM community Warstadt et al. (2023); Hu et al. (2024); Charpentier et al. (2025) has demonstrated that Transformer LMs, trained on cognitively plausible amounts of data, can often acquire fairly complex syntactic structures, even without instruction-finetuning. They can also learn accurate word-level representations when trained with character-level tokenization (Bunzeck and Zarrieß, 2025; Goriely and Buttery, 2025a). This makes them ideal testbeds for the aforementioned issue: does the construction distribution found in CDS, which features a high proportion of questions and syntactic fragments, affect the acquisition of formal linguistic capabilities? In other words, does robust linguistic knowledge at the word and syntax level emerge when the training data is closer to the fragmented, “messy” input of human learners?
@@
The goals of this paper, then, are twofold: (1) we compile a novel German BabyLM training set, for which we conduct the first utterance-level construction analysis for German. We find that distributions align with findings for English and other languages. We then (2) create three 5M-token subsets with distinct constructional profiles, varying, e.g., the proportion of fragmentary and complex utterances, and train small, character-based and subword Llama models on them. We evaluate them with lexical, syntactic, and semantic minimal pairs Bunzeck et al. (2025); Mueller et al. (2020); He et al. (2025) to gauge the influence of different construction distributions on these levels of linguistic knowledge, and find that differences between grammatically complex training data and a developmentally plausible constructional distribution are fairly small. While certain syntactic phenomena are learned somewhat better from more complex sentences, lexical learning improves with more fragments and questions in the input. Most interestingly, input complexity only modulates the steepness of the resulting learning trajectories, but has no principal effect on the amount of input needed to kickstart learning.
@@ 2 Constructions in children’s input
Child-directed speech can be seen as a separate linguistic register and is the primary input that children encounter in their first years. On the phonetic level, it features slower speech and exaggerated intonation patterns, which infants prefer listening to Zangl and Mills (2007), while its vocabulary is mostly restricted to everyday topics and children’s immediate surroundings Snow and Ferguson (1977). Structurally, child-directed utterances are usually shorter and simpler than adult-directed ones Genovese et al. (2020) and feature high amounts of structural and lexical repetition Tal et al. (2024). Statistical properties of the input directly influence the children’s order of acquisition for syntactic patterns Huttenlocher et al. (2002); Ambridge et al. (2015), e.g., for relative clauses Diessel and Tomasello (2000); Brandt et al. (2008); Chen and Shirai (2015).
@@
Early studies were mostly concerned with mapping out how much CDS is ungrammatical or otherwise “wrong” (in the sense of hesitations, false starts, etc., cf. Pine, 1994), but the quantitative turn in linguistics Janda (2013) has enabled more holistic analyses. In a seminal study, Cameron-Faulkner et al. (2003) analyze utterance-level constructions in child-directed English via corpora of toyplay sessions featuring children and caregivers. They show that CDS features only few “canonical” SV(X)-utterances but abounds with questions, lexical fragments, or copula constructions. The reported construction distributions also hold for typologically different languages, e.g., Irish Cameron-Faulkner and Hickey (2011). These constructions and their real-world functions help children to quickly understand the functional side of language. However, the most common and repetitive utterances that English-speaking children hear represent a rather skewed sample of the presumed, underlying formal language system. Generativist approaches would argue that certain formal processes, like question formation from relative clauses, are not attainable from this kind of language, as the input never contains specific examples Chomsky (1980) (although Pullum and Scholz, 2002 find that the input frequently contains exactly such specific examples). They also partly emphasize the importance of statistical learning, e.g. for providing hypotheses about competing possible mental grammars constrained by innate, language-specific mechanisms (cf. Yang, 2004, also Ambridge and Lieven, 2011, 121f.). Constructivist approaches do not view language learning as such a re-construction of the target language’s abstract grammar, but rather as the re-construction of the target language’s inventory of form-meaning pairings Behrens (2021). They argue that this kind of input is actually conducive to formal aspects of acquisition, by providing anchor points for first words and their semantic links to real-world reference, which then serve as building blocks for a gradual development into larger schemas (like questions with relative clauses).
@@
Although CDS features such a skewed construction distribution, written language aimed at children, e.g., in children’s books, is characterized by a much higher rate of canonical SV(X)-constructions than CDS Cameron-Faulkner and Noble (2013). Questions rarely occur in books. CDS produced in shared book reading presents a middle-ground — it contains more complex and SV(X)-constructions than regular CDS, but less than book text alone Noble et al. (2018). They argue that shared reading therefore, plays an important role in moving children from early, isolated traces of linguistic knowledge to a rich mental language system. This also aligns with the findings by Bunzeck and Diessel (2024), who show that the distribution of constructions in CDS varies with situation type (toyplay features most questions, meal sessions beget more imperatives, shared book reading features more complex constructions) and child age (questions and imperatives become less frequent with age). They suggest that CDS is therefore adapted to support children’s cognitive and linguistic development. Yet, as corpus studies are necessarily descriptive and cannot establish causal/mechanistic connections on their own (e.g. what would happen if a child never hears CDS), it remains questionable if this is actually true. Here, the potential of LMs trained on little data becomes apparent for constructionist approaches: they allow controlled experiments with different kinds of input data, which can serve as additional evidence for effects hypothesized from corpus data.
@@ 3 Input in developmentally plausible LMs
Early approaches to modeling language acquisition with neural networks used hand-picked, manually ordered data points Rumelhart and McClelland (1986) or synthetic data generated with hand-crafted grammars Elman (1993); Christiansen and Chater (1999); Chang et al. (2006). Both lack developmental plausibility. Since then, data availability has improved with the establishment of developmental corpora. Frequently, CDS from CHILDES MacWhinney (2000) is used to train developmentally plausible LMs (cf. Pannitto and Herbelot, 2020; Huebner et al., 2021). While CHILDES-based models have the advantage of learning from authentic data only, they have the disadvantage of not accessing the full breadth of the linguistic input children receive. Children are exposed to many more different registers of language throughout their linguistic development, like shared (or solitary) book reading, or television shows Montag (2019); Gowenlock et al. (2024). In response to this, the BabyLM corpora propose a data mix of varied spoken and written sources, from CDS over adult-adult conversations to OpenSubtitles Lison and Tiedemann (2016), but also children’s Hill et al. (2015) and adults’ books Gerlach and Font-Clos (2020). All data included in them could be plausibly encountered by children, which provides opportunities to ablate the influence of architecture/training on the learned linguistic knowledge.
@@
For languages other than English, data availability is the greatest problem for the construction of developmentally plausible datasets. Salhan et al. (2024) and Padovani et al. (2025) use only data available from CHILDES for models in different languages, whereas Prévot et al. (2024) compare models trained on spoken data (child-directed + adult-adult conversations) with models trained on the French Wikipedia. As such, these first forays into more polyglot BabyLMs are still constrained to the child-directed input found in CHILDES and do not extend to the aforementioned variety of inputs Soderstrom (2007); Gowenlock et al. (2024). Notably, Suozzi et al. (2025) introduce an Italian BabyLM but do not elaborate on their data sources beyond CHILDES.
Linguistic properties
The linguistic make-up of pre-training data and its influence on linguistic performance have only recently begun to receive increased scrutiny. Focusing on the lexical level, Yam and Paek (2024) measure sentence-level textual complexity with readability metrics based on text-wide word/syllable–sentence ratios for different corpora (CHILDES, BabyLM corpus, synthetic data, Project Gutenberg). They find that models trained on more complex text perform better at syntactic benchmarks, but simpler data (CHILDES) is learned better in terms of perplexity and loss convergence. Muckatira et al. (2024) filter English pre-training corpora for text spans that only contain vocabulary also found in English CHILDES data and find that simplified models generate more coherent text than models trained on more complex data, and also succeed in syntactic tests if the test data is filtered accordingly. In contrast, Edman et al. (2024) change the semantic content of the pre-training data and use datasets that approximate the linguistic input second-language learners get, e.g., dictionary entries, grammar books, and paraphrases. While grammar books moderately improve syntactic evaluation, there is no positive effect for the addition of the other text types.
@@
While actual research on the syntactic properties of the input is rather rare, training on filtered corpora has been used in pilot studies. Patil et al. (2024) and Misra and Mahowald (2024) filter out specific grammatical constructions from the BabyLM corpora and then probe the resulting models for knowledge of these grammatical constructions (which might also be analogically learned from related constructions or constructed from their parts). Patil et al. (2024) show that their models succeed on the BLiMP benchmark Warstadt et al. (2020), even if sentences containing structures targeted in BLiMP’s minimal pair sets are removed. Similarly, Misra and Mahowald (2024) show that acceptability scores for the English AANN construction can be reliably estimated from models that have never seen it. In sum, then, models appear to be able to generalize from indirect evidence and learn language in a somewhat constructivist, bottom-up fashion.
The structural composition of child-directed data has (so far) not been scrutinized. Most studies focus on lexical or semantic properties, emphasizing content over structure; child-directed data is usually equated with a somewhat fitting vocabulary or with just being authentic data. However, findings from usage-based linguistics suggest that structural properties, like utterance-level construction distributions, play a crucial role in language acquisition. Whereas Patil et al. (2024) and Misra and Mahowald (2024) remove specific constructions from their data, we aim to explore whether different global distributions of constructions influence the resulting linguistic knowledge and learning trajectories.
@@ 4 A German BabyLM dataset
To construct a German dataset, we use a variety of developmentally plausible sources, similar to the English BabyLM data Warstadt et al. (2023); Choshen et al. (2024). We use (1) all data from German CHILDES corpora MacWhinney (2000), including frog stories from TalkBank Berman and Slobin (1994) and math lessons from ClassBank Stigler et al. (2000), (2) subtitles from OpenSubtitles Lison and Tiedemann (2016), (3) adult conversations from the CallHome corpus Karins et al. (1997), and (4) written data from Project Gutenberg, from which we downloaded a manually curated sample of children’s books, young adult literature and literature commonly read in German schools. We supply this data with two corpora, the DreamBank database of self-reported dreams Domhoff and Schneider (2008) and short news texts from the Leipzig corpus Goldhahn et al. (2012); although they are not child-directed per se, these sources are child-available in everyday language.
To approximate child-available input even better, we tap into freely available child/learner-directed sources and compile four additional subcorpora for our dataset. The Wikibooks Wikijunior shelve features educational resources aimed at children, focusing on a diverse array of topics such as technology or nature. The Klexikon is a children’s wiki in German, featuring more than 3,000 articles aimed at children between 5–15. A simplified version of it is the MiniKlexikon, which features over 1,500 articles aimed at beginning readers. Finally, we also scrape the complete archives of Fluter, a magazine aimed at young adults published by the Federal Agency for Civic Education, which contains a large body of non-fiction. All resources are CC-licensed. Table 1 shows the raw token numbers for all corpora (16.5M overall). We extensively clean and normalize our data (details in Appendix B) and make our dataset available on Hugging Face. [^1] https://huggingface.co/datasets/bbunzeck/babylm-german